### v1

# Need Help? Designing Proactive AI Assistants for Programming

Abstract
While current chat-based AI assistants primarily operate reactively, responding only when prompted by users, there is significant potential for these systems to proactively assist in tasks without explicit invocation, enabling a mixed-initiative interaction. This work explores the design and implementation of proactive AI assistants powered by large language models. We first outline the key design considerations for building effective proactive assistants. As a case study, we propose a proactive chat-based programming assistant that automatically provides suggestions and facilitates their integration into the programmer’s code. The programming context provides a shared workspace enabling the assistant to offer more relevant suggestions. We conducted a randomized experimental study examining the impact of various design elements of the proactive assistant on programmer productivity and user experience. Our findings reveal significant benefits of incorporating proactive chat assistants into coding environments and uncover important nuances that influence their usage and effectiveness.
Figure 1: The implementation of a proactive chat assistant for programming. We introduce a proactive assistant that takes in the programmer’s context, which includes the current code and user message history, generates a set of suggestions, which include a summary, implementation, and explanation of implementation, and then determines whether it is timely to show the user in the interactive interface.
1 Introduction
Chat-based AI assistants such as ChatGPT [1] or Claude [2], enable people to accomplish some of their tasks via prompting: the user crafts a message with sufficient context about their task and then receives an AI response, they continue iterating this process until they are satisfied with the result. Common use cases of chat-based AI assistance include writing tasks [3, 4], idea generation [5], programming tasks [6, 7], and web navigation among other tasks [8, 9]. Interacting with these assistants requires two points of effort from the human: first, porting over the workspace context, e.g., copying the document or code file, and second, describing the task in natural language. Assuming the assistant had access to the user’s working context, a natural question is whether it could infer how it can help the user with the task they are trying to solve. If the assistant can reliably infer the user’s intent, generating a response to the task and suggesting it automatically might be advantageous. The assistant would now be proactive and result in a mixed-initiative interaction [10]. In this paper, we attempt to design effective proactive chat-based assistants powered by large language models (LLMs) for programming.
Virtual assistants have been explored in many commercial products including Apple Siri and Amazon Alexa [11]; in fact, one of the most well-known and first proactive virtual assistants is Microsoft’s Office Assistant, with the English version commonly referred to as Clippy. The Office Assistant would proactively offer assistance based on inference of the user’s intent from their Word document. While the user experience with Clippy was less than ideal [12], more recent examples of proactive assistants in the form of autocompletion for text-based tasks have been more successful. For instance, GitHub Copilot [13] provides in-line code suggestions with LLMs. Copilot has been well received by programmers and randomized studies have shown increased programmer productivity [14, 15]. However, autocomplete-based proactive assistance only considers local suggestions based on the user’s cursor position. In contrast, a chat-based proactive assistant could suggest modifications to or point out bugs in earlier parts of the code, in addition to being able to complete the code. Moreover, the particular design details of a proactive assistant can significantly impact its effectiveness.
We begin by outlining design considerations for designing chat-based proactive assistants with LLMs to address challenges of proactivity [16] by building on previous guidelines for human-AI interaction and mixed-initiative interaction [10, 17, 4]. Key considerations include supporting efficient evaluation and utilization of proactive suggestions and timing suggestions based on context. Given the success of LLM-based assistants for programming, the programming domain presents a good opportunity to study proactivity. Since the context of the domain is well-defined—consisting of the code file(s) in a programmer’s integrated development environment (IDE), integrating the proactive assistant into the programmer’s IDE enables us to overcome the challenge of observing the human workspace. Based on these design considerations, we build a proactive assistant for programming and integrate it into the chat interface within a code editor, as shown in Figure 1. The chat interface operates like a standard interface where users can type messages and receive responses; however, the assistant can now send proactive messages without a corresponding user request. The timing of the proactive suggestions depends on the human’s recent activities in the editor and their interactions with the assistant responses. Proactive messages are distinguished from regular assistant messages, with the former only displaying a high-level description unless the user chooses to expand the message to see a suggested implementation or explanation. Finally, each assistant message can be integrated into the code via a “preview” button highlighting the changes the message will make.
We conduct a user study to evaluate programmer usage of our proactive assistant against a baseline ChatGPT-like assistant. For our experiments, we use the web-based code editor RealHumanEval [18] as it allows us to easily conduct a controlled study to compare baseline and proactive conditions. To understand the effects of changing different design decisions, we investigate multiple variants of our proactive assistants, which differ in the timing of the suggestions and the ability to integrate suggestions into the programmer’s code. Our study demonstrates the benefits of proactivity in chat assistants for code as proactive assistants increase the number of tasks completed by an average of 12-18%. However, our study also shows that proactive assistants should be designed carefully as small changes can lead to differing user experiences. For example, 80-90% participants preferred the proactive assistant over the baseline chat assistant but only half the participants preferred proactive assistants that provided more frequent suggestions, despite productivity gains. We examined the telemetry and found that providing too many suggestions can saturate the utility of suggestions to users, leading them to copy suggested code into their editor less often. Furthermore, we observed that people tend to rely on proactive assistant capabilities when available, including the ability to incorporate code manually, rather than their manual efforts. Our findings demonstrate the potential of proactive assistants to improve programmer productivity and experience and highlight important design decisions that impact these downstream metrics.
In summary, this paper contributes:
  • A proactive chat assistant system, comprising an interactive interface, suggestion content based on the user’s code, and logic to appropriately time suggestions (Section 4). We also outline a set of design considerations when building proactive assistants (Section 3 and Section 7.1).
  • A controlled user study characterizing the benefits of our proactive chat assistant on downstream user metrics—e.g., productivity and user experience (Section 6.1 and Section 6.2), highlighting the potential for proactivity to be included in future coding assistants, as well as demonstrating the potential negative effects of changing certain proactive features. We also demonstrate how changes in proactive assistant capabilities lead to different interaction patterns. (Section 6.3).
2 Related Work
2.1 The Ecosystem of Programming Assistants
A growing number of tools powered by large language models (LLMs)—i.e., programming assistants—are available to developers to generate or edit code and answer queries. Programmers are increasingly writing code with AI assistants like Github Copilot [13] and Cursor [19] and are using chat assistants like ChatGPT [1] or Claude [2] in place of online Q&A communities like Stack Overflow [6]. Programming assistants typically involve one of two types of support: autocomplete suggestions are used to quickly write more code based on the programmer’s current code context, while chat dialogue can help answer questions that range from debugging errors to explaining documentation. Programming assistants surrounding code completion have been the focus of prior studies, which provide insight into how people use LLM-generated code completions [14, 7, 20, 21, 15, 22] and how to improve completions by providing explanations [23] and determining when to best show [24, 25]. While there has been less focus on investigating the use of chat assistants for programming, it is important to note that they complement assistance provided by code completions [26] and are the focus of this work.
Prior studies investigating chat dialogue for programming assistance have typically considered the setting where people initiate questions to the assistant [27, 28, 29, 6, 30, 31]. In particular, Chopra et al. [28] found significant time is spent constructing prompts and expressing their context to ChatGPT. As such, people did not enjoy being “slowed down” through this process. Additionally, Nam et al. [30] noted that people often had difficulty finding a good prompt for the chat assistant to provide the desired response. Further, Mozannar et al. [18] highlighted how the burden remains for programmers to incorporate an assistant response into their code. These observations highlight the hurdles to a streamlined integration of chat assistants like ChatGPT in a developer’s day-to-day workflow: the need to provide sufficient context about the developer’s problem and for developers to manually input and decide what they want to ask and then act upon the assistant response. Even while some programming assistants (e.g., Cursor [19], Github Copilot [13]) have begun building chat assistants into the integrated development environment (IDE), they are still reactive. In this work, we study whether incorporating proactivity into chat assistants can improve productivity and user experience. Our work aligns with existing tools for programming that provide chat support in that we still allow developers to ask questions when they so choose. Different from prior work, our proactive chat assistant periodically recommends suggestions and allows developers to integrate those suggestions into their code.
2.2 Background on Proactive Assistants
Research on proactive assistants has appeared in many forms, ranging from physical robots [32, 33, 34] to virtual chat assistants on phones or computer applications [35, 36, 37, 38]. Given our target application of programming, we focus our discussion on virtual chat assistants. Even within proactive chat assistants, the goals of prior proactive assistants vary depending on the context they are used. For example, support-based assistants use proactivity to provide motivation and continued dialogue [36, 38], while education-related assistants can proactively offer relevant support or explanations for students [39]. Most relevant to our domain are the set of information-finding assistants that leverage proactive messaging to provide additional and useful information [40, 41] and planning assistants can proactively help people reason about their decisions and identify overlooked alternative decisions or rationales [37]. The goal of our proactive assistant is to improve a programmer’s productivity while maintaining a good user experience; it requires blending multiple goals of existing proactive assistants. In Section 3, we discuss further how these downstream goals inform the design considerations for our proactive assistant.
Many prior deployed proactive assistants have been faced with significant negative reaction because the actual capabilities of these systems do not meet user expectations [11, 42], as many of the systems relied on pre-set messaging or simple models [35, 36]. With the advent of modern LLMs, and their growing usage as the backbone of agents [43, 44, 45], there is significant potential to revisit how we design more capable proactive assistants. In recent work, ComPeer [38] leveraged GPT-4 to build a proactive assistant for mental health support that handles and acts on a longer term memory of user dialogue. Aside from mental health applications, we believe LLMs can be particularly well suited for proactive assistants for productivity goals, particularly when they are already being used as regular chat assistants (e.g., ChatGPT). The increased model capability and ability to handle long contexts would be suitable for capturing complex environments in which they are completing tasks, which has been relatively unexplored in prior chat assistant literature relating to productivity [16]. In Section 4, we discuss the design of our proactive system which leverages off-the-shelf state-of-the-art LLMs to propose suggestions based on the current code context and even incorporate these suggestions into user code.
3 Design Considerations for Proactive Coding Assistants
To formulate design considerations for the core functionality and behavior of the proactive assistant, we revisit existing guidelines on designing mixed-initiative user interfaces and human-AI interactions [10, 17] as well as a prominent survey of known challenges of designing proactive chat assistants [16]. We highlight five design considerations that merit careful consideration to capture the benefits and challenges of proactivity.
Harnessing the benefits of proactivity. Horvitz [10] emphasizes the importance of “developing significant value-added automation” when developing effective mixed-initiative systems. Proactivity can address the limitations faced by programmers when interacting with existing, reactive chat assistants for coding contexts. As discussed in Section 2.1, the value-add of introducing proactive suggestions would be an improvement in programmer productivity (e.g., in terms of time and effort) and in the user experience (e.g., reducing perceived “slow-down”), as compared to a non-proactive chat assistant [28, 30, 18]. To ensure that programmers can effectively use proactive suggestions, we propose the following two design considerations:
(1) Support efficient evaluation. Proactive suggestions should be provided in a way such that programmers can efficiently interact with them. This involves providing a manageable amount of information to the user about the suggestions so that they can quickly decide if they are worthwhile. This might involve hierarchically scaffolding the suggestions so the programmer can obtain more information about the suggestion if needed.
(2) Support efficient utilization. If the user decides the suggestion is useful, the proactive assistant should make it easy for the programmer to utilize. For example, the proactive assistant should make it as easy as possible for them to take action corresponding to the suggestion and incorporate it into the programmer’s code. If the programmer decides the suggestion is not useful, they should be able to dismiss the assistant easily.
Handling challenges of proactivity. Prior works on proactive chat assistants have demonstrated that there is considerable nuance in developing a proactive chat assistant [16], as irrelevant and untimely proactive messages may compromise the success of the interaction timing [46, 35, 47, 48]. We incorporate these challenges in our design considerations:
(3) Show contextually relevant suggestions. Proactive chat assistants are increasingly built into complex environments where users often perform multiple different tasks. For example, programming tasks can range from implementation to debugging and testing, to name a few, that require different interventions and information from the assistant. Proactive assistants should make sure to surface information that is relevant to the current context and recent user interactions and messages.
(4) Time suggestions based on context. Programmers perform various tasks during their work and are often in a state of flow [49]. The proactive assistant should consider what the programmer is currently doing and their workflow before it provides a suggestion [20]. For example, showing suggestions while the user is actively coding can disrupt the programmer’s flow and lead to disruptions. Horvitz [10] discusses the importance of fallback behavior which can allow for a reduced frequency of suggestions.
(5) Incorporating user feedback. The proactive assistant should adapt its suggestions and timing as it interacts with the user. The user should be able to accept and reject proactive suggestions, and the decision to accept or reject suggestions should influence future decisions of the proactive assistant. This is especially important as different programmers may want different levels of proactive help from the assistant.
In the following section, we introduce an implementation of a proactive assistant for programming building on the design considerations above.
4 Proactive Assistant Implementation
In this paper, we build a proactive assistant for programming tasks based on the design considerations in Section 3. Our proactive chat assistant distinguishes itself from online AI assistants such as ChatGPT in two ways: first, our proactive assistant has access to the user’s work context, and second, our proactive assistant proposes suggestions instead of passively waiting for user requests. When embedding the proactive assistant into the user’s work context—i.e., the programmer’s IDE, the proactive assistant will have access to both the current code in any of the user’s files and prior user queries to the AI assistant. To turn context into suggestions, we discuss two primary components that comprise our proactive assistant, as shown in Figure 1, and how they instantiate the set of design considerations:
  • Interactive interface: In Section 4.1, we overview the chat interface which shows suggestions from the proactive assistant in a way that is easy for users to interact with and understand (Design Considerations 1-2) and to provide feedback (Design Consideration 5).
  • Suggestion generation: In Section 4.2, we discuss how the assistant takes the current context and generates a set of proactive suggestions that is timely and relevant (Design Consideration 3). In Section 4.3, we discuss how to appropriately determine when suggestions are shown to users (Design Consideration 4).
4.1 Interactive Interface
Figure 2: A walkthrough of the proactive assistant interface for coding. (A) Overview of suggestions, where the assistant provides a short description, and users can choose to expand a suggestion for more details (this can involve implementation and a brief explanation); (B) Buttons that allow participants to preview the implementation in their code, accept a suggestion to ask follow-up questions on, or delete the suggestion; (C) Integrating a suggestion into the editor via a diff format, where users can decide if they want to accept or hide the changes; (D) Invoking suggestions, where users can also request suggestions from the assistant (for example, if they did not like the original suggestion or wanted a suggestion before the assistant provided one).
The interface of the proactive assistant, shown in Figure 2, is built into the standard chat interface, allowing programmers to use all normal chat functionality, while the assistant periodically provides suggestions to programmers. We describe the different features of the proactive suggestion and their respective design considerations:
4.1.1 Suggestion summary.
Towards fulfilling design consideration 1, the proactive assistant facilitates efficient evaluation of the suggestion by providing a summary to allow users to decide whether the suggestion is relevant. The summary consists of a single sentence that starts with the type of the suggestion e.g., a bug fix, a new feature, and then a description of the suggestion, as shown in the headings of Figure 2 (A). The drop-down style of the interface allows users to browse the various suggestions easily. Through pilot studies, we find that showing a set of three suggestions strikes the balance between creating a diverse yet manageable set of information for users to look at. We explore the effect of changing this parameter in our user study.
4.1.2 Details of suggested implementation.
If the user decides to look into a suggestion further they can expand it for more details as shown in Figure 2 (A). Depending on the type of suggestion, they may receive a code snippet and/or a text description (e.g., a suggestion that explains a code snippet may not include code). The interface also allows users to copy the suggestion into the editor (design consideration 2). We limit the text description of the implementation to a few bullet points to facilitate efficient interaction (design consideration 1).
4.1.3 Preview suggested implementation in code.
While users can copy the suggestion and decide how to use it themselves, the proactive assistant also has a “preview” functionality where the assistant will suggest an option for how the suggestion can be incorporated into the code (design consideration 2). The suggested implementation is computed via another call to the LLM, where it is given the proactive suggestion and the user’s code and produces new code incorporating the suggestion. This generated code file is then shown to the user via a diff editor based on the Monaco Editor [^1] https://microsoft.github.io/monaco-editor/ which tells the user what lines were added or removed once the suggestion was incorporated, as shown in Figure 2 (C). Once presented with the preview, the user decides whether to accept or hide the changes. They also accept a subset of the changes by editing freely before clicking on the “accept changes” button. In the Appendix, we include the prompt that incorporates edits into the current code snippet.
4.1.4 Asking follow-up questions based on suggestion.
A natural interaction paradigm when using chat assistants is for users to provide clarifications or ask follow-up questions based on the assistant responses. Since the proactive assistant is built into the chat interface, we tap into this paradigm when scaffolding interactions for proactive suggestions. As such, we create the option for users to add a proactive suggestion to the chat history (via the “accept” button) or remove it (via the “delete” button). Clicking on the accept button shown in Figure 2 (B) would add the proactive suggestion to the chat message in a way that looks visually similar to the rest of the user’s messages. This also allows the user to ask follow-up questions on the proactive suggestion. This may be helpful as it is common for users to follow up on LLM responses to clarify or correct the model. Further, accepting and deleting suggestions allows future proactive suggestions to incorporate user feedback, addressing design consideration 5.
4.1.5 Manually requesting suggestions.
While the nature of a proactive assistant suggests that the assistant will determine when to provide suggestions to the user, it is unrealistic to expect an assistant to anticipate when users might want suggestions (design consideration 4). As such, we include a button, as shown in Figure 2 (D), to allow users to request suggestions. This can also allow us, in future work, to better learn when the proactive assistant should intervene.
4.2 Suggestion Content
Suggestion Type Example of developer question
Explaining existing code What does this do?
model = GPTLanguageModel()
m = model.to(device)
Brainstorming new
functionality
Based on the following OCaml code for an s-expression evaluator, write a parser for the tokens defined. Note that all operations must be binary and bracketed, with no concept of operator precedence.
[code context]
Completing unfinished code The following is some C code for binding a "hello" C function to an Io method. Can you complete the code by the comment:
[code context]
Pointers to documentation I’ve saved this in ‘scripts/align-import.py’.
Remind me how to use ‘find’ to run it on every ‘*.lean‘ file in a subdirectory?
Debugging
(Latent errors)
Why is my redirect not working? Here is my client side code
[code context]
Everything else works as intended, except that it will not redirect. What is the issue here?
Debugging
(Runtime errors)
[code context]
Test suite failed to run
Cannot find module ‘libp2p’ from ‘src/shared/libp2p_node/index.ts’
[remainder of error trace]
Do you have any idea why I am getting this error?
Adding unit tests Could you create Jest unit tests for this function?
[code context]
Improving efficiency and
modularity
The following is a kernel of an algorithm. It uses Apple’s metal api for matrix operation. I think it can be improved to make it run faster.
[code context]
Table 1: Suggestion types informed by dataset analysis. We identify categories of suggestions that a proactive assistant can surface by coding developer questions to ChatGPT in the DevGPT [6] dataset. We incorporate these suggestion types into the design of our proactive assistant.
For the proactive assistant to be useful, it must generate appropriate suggestion content (design consideration 3). This means that the assistant should understand what types of suggestions will benefit users and that the assistant should surface the right type of suggestion at the right time, depending on the user’s code context.
To identify the different types of suggestions that users might ask and benefit from, we analyze DevGPT [6], a dataset of conversations between programmers and ChatGPT, as a source of real-world questions. While other chat datasets also contain in-the-wild programming-related questions (e.g., WildChat [50], LMSYS-Chat-1M [51]), we selected DevGPT in particular because the majority or all questions contain user code context in addition to the user question and LLM response, which allows us to manually inspect whether it would be reasonable for a proactive assistant to anticipate this question from the current code context. We sample 100 questions and their code contexts from the DevGPT data and manually create a label for the “category” of the question. We then cluster labels to identify common themes across questions. The goal of surveying existing questions is to ensure that our proactive assistant can provide diverse suggestions; however, the categories we identify may not be exhaustive.
From this annotation process, we identify eight categories of questions that proactive assistants can help with: explaining existing code, brainstorming new ideas or functionality, completing unfinished code, providing pointers to syntax hints or external documentation, identifying and fixing bugs (which include both latent and runtime errors), adding unit tests, and improving code efficiency and modularity. In Table 1, we include example user questions that fall under each category. To scaffold these suggestion types into proactive suggestions, we incorporate them into the system prompt to encourage the model to consider a more diverse set of suggestions. In the Appendix, we include the prompt used to generate the proactive suggestions.
4.3 Timing of Suggestions
The proactive assistant must time suggestions appropriately to not be obstructive or distracting to users (design consideration 4). A study by Barke et al. [20] characterized interactions with programming assistants as consisting of two modes: acceleration, where the assistant is primarily used to help the programmer implement what they already intend to write; and exploration, where the programming assistant assists a programmer in identifying and planning out goals. We propose the following conditions for when proactive suggestions should be shown to users based on an estimation of when the user is in acceleration or exploration mode:
  • During acceleration: Suggestions are not requested when the user is interacting with a previous suggestion, typing in the chat, sending a message, or waiting for a response. The suggestion timer resumes 5 seconds after the user stops typing to account for a typing break. If the user would like a suggestion, they can manually request one.
  • During exploration: When the user is idle—likely due to planning and brainstorming, the assistant provides suggestions every 5 seconds, limited to every 20 seconds since the last suggestion, interaction, or chat. If the user starts coding before suggestions are shown, due to potential latency in the model query, we do not display the suggestions.
Additionally, we show a proactive suggestion during debugging, when users are trying to edit the current code for any bugs or improve the performance. Debugging is distinct from the previous two states because the user is neither trying to write more code nor planning the next steps. The proactive assistant is immediately triggered when the user runs or submits code that leads to an error because the user is not actively coding at the moment. We modify the suggestion generation prompt to incorporate the error statement and restrict the types of suggestions that the assistant can generate. In the Appendix, we include the prompt used to generate suggestions during debugging.
5 Study Design
We conducted a controlled user study to understand the effectiveness of a proactive assistant on productivity and investigate how people interact with the assistant. Our research questions are as follows:
  • RQ1:
    What is the effect of the proactive assistant on user productivity? If the proactive chat assistant can anticipate user needs or even surface issues the user has not considered, we expect the assistant to have a measurable downstream effect on productivity measures.
  • RQ2:
    What is the effect of the proactive assistant on user experience? In addition to productivity measures, the user should have a favorable experience when interacting with proactive assistants. If the proactive assistant is providing sufficient perceived value-add, the user would ideally perceive the proactive assistant to be more helpful than a standard assistant like ChatGPT.
  • RQ3:
    How do participants interact with different proactive assistants? A proactive assistant comprises several design considerations, we analyze how users interact with different variants of our proactive assistant to understand the effect of these design decisions on user behavior.
5.1 Participants
We recruited a total of 61 students via university mailing lists. The inclusion criteria for the study are that they must be based in the U.S., be older than 18 years old, and have experience programming. Since we design tasks that are written in Python, we require participants to have a baseline level of Python knowledge. Among our participants, 60% reported their Python knowledge as intermediate, 31% had advanced Python experience, and 9% were beginners. Moreover, 19% self-report that they are daily users of AI tools for programming (including GitHub Copilot and ChatGPT), 46% report that they use these tools at least once if not multiple times a week, and 35% rarely use AI tools for programming.
5.2 Baseline
Our study compares proactive assistants to a non-proactive chat-based assistant baseline. Every participant in our study interacts with a standard chat assistant that is incorporated into the coding environment but does not have access to the user’s code. This is similar to the set-up that has been studied in prior papers [31, 6, 18] and mimics how people would use ChatGPT. The interaction paradigm for the baseline is as follows: after the user types in their query, which may contain code from the IDE to provide relevant context, the chat assistant can respond in natural language and provide code snippets. To utilize the code snippets, the user would need to copy in the assistant response back into the IDE.
In both the baseline and our experimental conditions, we do not include any code autocomplete tools to not introduce a potential confounding factor in our analysis as not all participants are experienced in using autocomplete tools. Further, we believe the proactive assistant can complement code autocompletion tools, which consist of local proactive changes, while the proactive assistant provides suggestions at a more global level.
5.3 Experimental Conditions
We consider three experimental conditions to explore the effect of different proactive chat assistant designs:
Condition 1: Our proactive chat assistant (Suggest and Preview):
In this condition, the participant interacts with the proactive assistant described in Section 4. This condition encompasses all features of the baseline condition, i.e., the participant still has the option to send chat messages to the assistant at any point while also receiving proactive suggestions.
Condition 2: Our proactive chat assistant without preview feature (Suggest):
In this condition, the participant interacts with the proactive assistant described in Section 4, without the preview feature that allows the assistant to suggest direct edits to the user’s code. Similar to Condition 1, the participant can still send chat messages to the assistant at any point while also receiving proactive suggestions.
Condition 3: Persistent proactive chat assistant (Persistent Suggest):
In this condition, the proactive assistant in this condition looks at face value to be the same as the proactive assistant in Condition 2 (Suggest). However, we modified a few different parameters that were discussed in Section 4: we reduced the amount of time that an assistant waits to provide a suggestion (from 20 seconds to 5 seconds), increased the number of suggestions shown (from 3 to 5), and removing guiding prompts.
The participants were randomized and roughly divided equally across the three conditions, with 20 in Suggest and Preview, 21 in Suggest, and 20 in Persistent Suggest. Comparing the three conditions allows us to investigate whether (1) the level of proactivity affects user productivity and user experience (comparing Suggest and Preview and Suggest) and whether (2) small changes in the design of a proactive assistant can make a sizeable difference in user productivity and experience (comparing Suggest and Persistent Suggest). We fix the LLM backbone across all conditions to be GPT-4o [1], a state-of-the-art LLM.
5.4 Experimental Platform
The study was conducted online and asynchronously at the participant’s own time. The interface we used to deploy the study is an extension of the open-source platform RealHumanEval introduced in [18]. Conducting the study on the web allows for easy assignment of participants to experimental conditions, the duration of the study, and the tasks participants solve. Finally, the web interface reduces installation issues or any potential incompatibility. While the web interface is not the traditional IDE that a user may typically use for their day-to-day development, it is reminiscent of online coding platforms (e.g., Leetcode) that participants are generally familiar with. The editor in our web interface is also the same Monaco editor as in Visual Studio Code, a popular IDE. We also added a check in the post-study form to make sure that participants did not have any issues using the interface.
5.5 Procedure
Before participating in the study, each participant filled out a consent form. The study has been approved by our institution’s review board (IRB). We adopt a between-subject setup to compare the three proactive chat variants and a within-subjects setup to compare each proactive variant to the non-proactive baseline. This means each participant will interact with both a proactive assistant and the baseline chat condition. We randomize the order in which participants interact with either the proactive assistant or the baseline chat assistant. The total time the participants spent coding for the study was 40 minutes, with participants spending 20 minutes in each condition.
5.5.1 Onboarding
Before participants can access the proactive assistant, they are provided with a short tutorial that describes how to interact with the proactive suggestions. The tutorial highlights how the proactive assistant will provide a high-level summary of suggestions that can be expanded upon and how to ask follow-up questions. The tutorial also warns participants that the suggested implementations may not always be correct and to verify the suggestions carefully. In the Appendix, we provide a copy of the onboarding instructions that participants had access to.
5.5.2 Coding Tasks
In our study, participants engage in 20-minute sessions focusing on two types of programming problems designed to reflect real-world engineering scenarios. The first type of problem aims to test system-building skills. For instance, participants are tasked with enhancing an “online store” class object implementation. Given an initial starter code, they must implement additional features specified through natural language instructions, brainstorm a new feature, debug potential issues, and write tests for these new features. The second type of problem challenges participants to work with unfamiliar packages, simulating situations where engineers must quickly adapt to new tools. An example of this is manipulating a dataframe using various NumPy functions. Unlike previous studies that primarily assessed puzzle-solving abilities through Leetcode-style problems [14, 18], our approach aims to evaluate skills more directly relevant to practical software engineering tasks. To avoid confounding task difficulty or type with the helpfulness of proactive assistants, participants will start with the same task type in both proactive and baseline conditions, though the tasks themselves are randomized. In the Appendix, we provide descriptions for each task in the study.
5.5.3 Post-task survey
The study concluded with a post-task survey that asked participants about their experience with both the proactive and baseline conditions. While we log telemetry data in both conditions to measure user productivity, we use the post-task survey to measure user perceptions. In the Appendix, we provide a list of the post-task survey questions.
5.6 Measurements
To answer our research questions, we measured the following:
  1. 1.
    Telemetry. Our interface logs all user behavior from coding to using the chat assistant: anytime the code is updated by the user, the interface saves the updated code. Further, any chat messages and associated responses are logged. Finally, any interactions with the proactive assistant, including expanding suggestions, accepting a suggestion, and requesting suggestions are all recorded. Using the collected telemetry data, we compute the the number of sub-tasks completed. We can also obtain fine-grained interaction behavior to identify common interaction patterns with proactive suggestions.
  2. 2.
    Perceived usefulness. In the post-task survey, we ask participants to rate their interactions with both the proactive and baseline chat assistants. Participants were then asked to explain their ratings.
  3. 3.
    Open-ended feedback. In the post-task survey, we also asked additional questions that further dive into what aspects of the proactive assistant participants preferred and what could be improved.
5.7 Analysis Approach
We measure the effect of proactive assistants on user productivity (RQ1) by computing the number of sub-tasks completed across each of the four coding tasks. We build a linear model which incorporates the experimental condition, which is either baseline or one of the proactive models, and the coding problem as fixed effects. We also qualitatively inspect whether the code written contains automated test cases, which is supplementary measure of well-written code. To understand the effect of proactive assistants on user experience (RQ2), we measure participant ratings of the baseline and proactive assistant. We then computed a binary measure of whether participants preferred the proactive assistant to the baseline and ran an ANOVA with Tukey HSD. We also analyze participant’s written justifications to interpret the quantitative results. For all tests, the threshold for statistical significance was $\alpha=0.05$. To begin to understand how participants use the different proactive assistants (RQ3), we measure how often participants used the various components of the proactive assistants as logged in the telemetry (i.e., expanding, previewing, accepting, copying, and requesting suggestions). We consider this portion of the results an exploratory analysis to identify trends that distinguish the proactive conditions and understand the effect of different design decisions on user interactions. We present the results of our study in the following section.
6 Results
6.1 RQ1: Proactive Assistants Generally Improve User Productivity
Figure 3: Percentage of sub-tasks completed correctly. Comparing baseline chat to proactive assistants across the four tasks, where Task 1 and Task 2 are system-building questions and Task 3 and 4 are ones where participants work with new packages and functionality. While performance varied by task, we observed that all variants of proactive assistants tended to increase the number of test cases passed compared to the baseline chat assistant across the board.
Number of sub-tasks completed.
When comparing the number of sub-tasks completed by participants, we find that on average participants with proactive assistants are more productive with a baseline chat assistant: we observe a $12.1\%\pm 5.1\%$, $18\%\pm 5.8\%$, and $11.6\%\pm 5.0\%$ increase in the percentage of test cases passed for Suggest and Preview, Suggest, and Persistent Suggest respectively compared to baseline. Figure 3 provides a more granular view by task. We find that the improvements in the number of test cases over the baseline chat assistant are significant across all proactive variants, where $p=0.01$ for Suggest and Preview, $p=0.002$ for Suggest, and $p=0.02$ for Persistent Suggest.
Number of test cases written.
In addition to the increase in the number of test cases passed, we also observe an increase in the amount and quality of test cases written. Note that writing tests was part of the instructions provided to users. Focusing on Task Type B, where it is important to test how different components of the designed system work together, we annotate whether the user code contains test cases. We find that only 13.3% do in the baseline setting while 33.3% do in the proactive settings. We believe that providing proactive suggestions can help create a virtuous cycle where the suggestions generate tests for code the participant has written to surface issues in their code, thus increasing the number of test cases passed.
6.2 RQ2: User Experience with Proactive Assistants Varies by Implementation.
Figure 4: Comparing participant perception of proactive versus baseline. User experience is important to the adoption of proactive assistants. We compare how often participants preferred a proactive assistant to the baseline variant in each condition. We find participants generally preferred the proactive variant only in the Suggest and Suggest and Preview conditions.
User preferences across proactive conditions.
While we observe that proactive assistants generally improve user productivity, user experience and perceptions of the proactive assistant vary across the conditions. As shown in Figure 4, we measure whether participants preferred the proactive assistant over the baseline non-proactive assistant. We find that the vast majority of participants prefer the proactive variant over the baseline for both the Suggest and Suggest and Preview proactive assistants (90% and 80.1% of participants respectively). In contrast, less than half of participants (45%) in the Persistent Suggest proactive condition preferred having the proactive assistant. We find that both Suggest and Suggest and Preview variants are statistically different than the Persistent Suggest variant ($p=0.005$ and $p=0.03$ respectively). On average, participants had a neutral view of the baseline assistant in terms of usefulness. As such, participants viewed both the Suggest and Suggest and Preview variants on average as net beneficial.
Comparing participant responses across proactive variants.
Participants in the Persistent Suggest condition tended to not prefer the proactive assistant because they often found the assistant to be “distracting” and “annoying”. One participant noted that “the non-proactive chat assistant was best because it didn’t interrupt what I was doing.” and another participant “found it annoying because it distracted [them] from working with the non-proactive answers.” In contrast, participants in the Suggest and Suggest and Preview conditions tended to have a positive view of the proactive assistant. One participant noted that the reason they preferred the proactive assistant was because “non-proactive was pretty useless because it didn’t have any of the context regarding what I was trying to do or the task; I could only really use it for pure syntax like what I would usually search StackOverflow for. Proactive was better because it had more context.” and another stated that “I really wasn’t sure what to ask for with the non-proactive chat” since the baseline chat “required manual input to generate advice”. Next, we analyze how user perception translates into the actual usage of the proactive assistant.
6.3 RQ3: How Do Participants Use Different Proactive Assistants?
To better understand the effect of different design decisions on how participants use proactive assistants, we break down user interactions with proactive suggestions by condition. First, we compare Persistent Suggest and Suggest conditions to explore the effect of suggestion timing. Second, we compare Suggest and Suggest and Preview conditions to explore the effect of the level of proactivity.
Figure 5: Comparing interactions with different proactive conditions. Visualizing sampled participant trajectories from coding with different proactive assistants. We denote “expanding” a suggestion with E, “accepting” a suggestion with A, “previewing” a suggestion with P, and “copying” a suggestion with C. In the Persistent Suggestion condition, participants tend to expand suggestions without utilizing them further. In contrast, participants tend to copy expanded suggestions in the Suggest condition while participants in the Suggest and Preview condition tend to preview suggestions.
6.3.1 The effect of suggestion timing.
Decreasing the amount of time that the proactive assistant waits to provide suggestions means that far more suggestions were shown in the Persistent Suggest condition than in the Suggest condition. As such, we observe that participants tended to expand more suggestions in the Persistent Suggest condition compared to the Suggest condition ($9.5\pm 2.4$ times per task versus $6.4\pm 2.5$ times). However, proactive suggestions were copied more often in the Suggest condition compared to Persistent Suggest ($3.1\pm 1.5$ as compared to $2.1\pm 0.9$ times). Figure 5 shows example telemetries of participants in each condition that demonstrate this behavior. These trends suggest that when a proactive assistant provides too many suggestions, participants can be easily distracted and thus still want to view the suggestions at a similar frequency. However, when the suggestions are provided too often, participants are less likely to make use of the suggestions.
6.3.2 The effect of level of proactive assistant involvement.
The ability to preview and incorporate suggestions into code changed how users interacted with the respective proactive assistants. For example, after expanding a suggestion, further interactions with suggestions were more than twice as likely to be through the preview functionality rather than copying code and integrating it into the editor by themselves ($3.1\pm 1.5$ in the Suggest and Preview condition compared to $0.8\pm 0.5$ Suggest condition)—an example of such an interaction pattern is shown in Figure 5. This may also have led to an increase in manual requests for suggestions, which happened more often in the Suggest and Preview condition, where participants made on average $2.9\pm 0.9$ manual requests in the Suggest and Preview compared to an average $1.7\pm 0.7$ requests in the Suggest condition. These results suggest that participants may gravitate towards options that reduce their manual efforts to make changes, which include utilizing the preview functionality.
7 Discussion
7.1 Design Implications For Proactive Coding Assistants
7.1.1 Revisiting design considerations
Given the findings from our experimental study, we now revisit whether the design considerations we posed in Section 3 are useful guides for designing proactive coding assistants. We map qualitative feedback from participants’ post-task survey responses to the five design considerations.
(1) Support efficient evaluation. We observe that participants generally felt that the proactive assistant provided suggestions that had “clear answers while giving specific examples” and allowed them to quickly “get a baseline understanding of the problem.”, which provides evidence that the suggestions were presented in a way that allowed participants to understand and use them efficiently. Further, we believe the suggestions were not too distracting; another participant noted that the suggestions helped to “pinpoint specific issues while not being so obnoxious [they] couldn’t ignore it to focus on other tasks.”
(2) Support efficient utilization. While both the baseline non-proactive and proactive assistants can help “find errors in my code and write functions”, a participant noted that they are “able to work faster while using the one with suggestions.” In particular, many participants noted that proactive suggestions were “much easier and faster to just click on a roughly correct suggestion instead of typing out specifically what I wanted to know” and “a lot of the suggestions can be directly pasted into the code.”. Further, proactive suggestions that can be directly incorporated into user code via the preview function are “immediately executable”, and participants felt that it improved their general workflow efficiency.
(3) Show contextually relevant information. Not only did participants believe that the proactive assistant often suggested relevant information (e.g., “it gave me suggestions which were related to what I wanted”), but the assistant also often “suggested things I didn’t consider and didn’t know to ask for such as refactoring code and spelling errors”. Identifying errors seemed to be a common feature that participants liked, as “it took care of most of the errors and allowed me to focus mainly on writing my code”. Overall, participants felt the proactive suggestions were “more personalized” and “more tailored” to their needs. In particular, a participant contrasted the proactive chat assistant with other programming assistants such as code completion tools as “[code completion] only fills the code after my cursor and don’t provide insights on the already implemented code.” This reinforces the idea that proactive chat assistants can complement commonly used tools like GitHub Copilot.
(4) Time services based on context. Aside from the Persistent Suggest condition where participants felt suggestions appeared too often (as discussed in Section 6.2), participants had positive perceptions of the proactive suggestion timings, noting that the proactive assistant was “always ready” and comparing the proactive assistant to “a second set of eyes” or “a second person solving [the problem] for me by my side.” Additionally, multiple participants suggested that the timing of suggestions can even enable them to “catch issues early” or “fix errors early on so that they don’t require refactoring the entire codebase later”. Our results across proactive conditions demonstrate that appropriately timed suggestions can be very helpful to users, however inappropriately timed suggestions may be perceived as more of a nuisance.
(5) Incorporating user feedback. Interestingly, this was the only design consideration that received mixed feedback from participants. Some participants believed “the ability to accept or reject the code it suggested was really nice” and “intuitive” to use while others felt like they were “unnecessary” and “interacting with them distracts me from the task.” Future work may consider incorporating user feedback implicitly rather than requiring users to accept or delete suggestions.
7.1.2 Expanding design considerations
We also asked participants how the proactive assistant can be improved and synthesized these responses to identify three new design considerations. We do not include features that were already accounted for in the different variants (e.g., participants in the Persistent Suggest condition wanted less frequent suggestions, and participants in the Suggest condition wanted to be able to incorporate suggestions into the code). These expanded design considerations present opportunities for future work to improve proactive coding assistants.
(6) Allow users to decide when they want proactive assistance. Multiple participants across conditions recommended having an option for users to decide when to turn the proactive suggestions (e.g., “an opt-in feature” or “could change the frequency”), particularly because “sometimes I found it useful and sometimes I didn’t”. Since the goal of our study was to investigate the usage patterns and effects of a proactive assistant, providing the user the option to turn off the proactive assistant would not have been possible. However, this feature could be important for facilitating appropriate long-term usage, as a participant was concerned about potential over-reliance on proactive suggestions, which “can lead to a lack of understanding of a codebase. Therefore, I don’t want constant suggestions.”
(7) Incorporate additional context when generating suggestions beyond user code. While many participants already appreciated the benefits of proactive suggestions, they had further ideas for improving the suggestion content by incorporating other aspects of user behavior. For example, participants suggested that the proactive assistant could take into account “the functions I was using’’, “what the user is focusing on, typing, or clicking,” or “recently updated code regions.” Another participant suggested that it may be beneficial for the user to incorporate further context by providing a natural language of their task—i.e., “if I can enter the task description somewhere as a constant part of the prompt [as] some suggestions have a slight misunderstanding of the problem”. These recommendations demonstrate that augmenting the existing code context with additional user information may be able to further tailor suggestions to user needs.
(8) Flexibly varying the amount of information shown. Since the current proactive assistant provides one implementation option and a fixed number of suggestions each time, participants noted the possibility of scaling the amount of information up and down. For example, the assistant could provide more implementation options in the proactive suggestions, e.g., “multiple ways of solving a problem,” particularly when different factors may be important (e.g., time versus space efficiency). However, there may also be times when the assistant could scale down the number of suggestions shown (e.g., “if everything was good, then tell me that everything is good.”)
7.2 Limitations
There are a few limitations to our study results and setup which we discuss here. First, our study is time-limited and consisted of tasks that were given to the participants rather than the tasks being self-motivated by the participants. The coding tasks we relied on do not span the entire set of tasks a professional programmer might encounter in their work. The tasks were also easily measured in the number of test cases passed, which may not be the best metric in practice. Second, the participants in our study were entirely comprised of undergraduate and graduate students, which is not fully representative of the audience that uses AI tools for programming. Third, our implementation of the proactive assistant was integrated into a web-based coding environment which does not contain the full features of an enterprise IDE. Fourth, the conditions that we considered in our controlled lab study were limited to chat-based systems, since not all participants had experience using GitHub Copilot.
Further work is necessary to understand the nuances of proactive assistants on a wider range of user backgrounds and coding tasks. Future work could consider evaluating the proactive chat assistants in the presence of other programming tools like code completion to understand the interplay between different tools and their relative effects on productivity.
Appendix A User Study Details
A.1 Instructions
General instructions
Welcome to python coding study!
  • You will be writing code in Python only, and use only standard python libraries in addition to numpy.
  • You will have 40 minutes total where you will try to solve as many coding tasks as possible one at a time.
  • It is NOT allowed to use any outside resources to solve the coding questions (e.g. Google, StackOverflow, ChatGPT), your compensation is tied to effort only.
Baseline condition
You will write code in the interface above: a code editor equipped with an AI assistant chatbot. The chatbot does not have access to either the code editor or the task description.
  • Please write python code only in the editor. We only support standard python3.8 packages and numpy.
  • You can run your code by pressing the "Run" button and the output will be in the output box at the bottom in grey.
  • Press the “Submit” button to submit your code. You can only submit your code once for each task.
  • You are provided with a chat interface that you can use to help you with the coding tasks.
  • Pressing the "Clear" button on the top of the chat window will clear the current chat history. Chat history will be automatically cleared when you move to the next task.
  • If the chat output contains a code snippet, you can click the icon to copy the code to your clipboard. Please be aware that its output is not always correct.
Proactive condition
You have access to the Proactive Coding Assistant. In addition to the regular chatbot functionality, the assistant will occasionally provide suggestions when you are stuck! Please try to use it in the study if the suggestions seem helpful. Here’s how:
Click on a suggestion title for the following options:
  • Previewing: Click "Preview" to see how the proactive assistant incorporates the suggestion into your code, which can then accept or hide the changes. Use the copy icon to copy the code to your clipboard.
  • Accepting: click “Accept” to add the expanded suggestion to the chat window and ask follow-up questions.
  • Deleting: click “Clear all” to remove all current suggestions. Use “Delete” to remove a single expanded suggestion.
Other features:
  • If a regular chat message contains code, you can also have the proactive assistant help you incorporate into your code.
  • You can request suggestions by clicking on the “Suggest” button at the top of the chat window, or using the shortcut [Ctrl + Enter] (Windows) or [Cmd + Enter] (Mac) in the code editor.
  • After running your code, the assistant analyzes the output or error messages and offers debugging suggestions. You can interact with these suggestions in the same way.
Warning: the assistant does not have access to the task description, but it has access to your code editor and can provide context-aware suggestions based on your code and cursor position. Suggestions are not always perfect, and the code provided may be inaccurate. In addition the suggestions may use packages that we do not support (we only support numpy). Always review suggestions thoroughly before integrating them into your code.
You can review this tutorial again by clicking on the “Show Instructions” button at the top of the page, then clicking "Next".
A.2 Post-study Questionnaire
  • Overall, how useful was the non-proactive chat assistant (e.g., the one that did not automatically provide suggestions)? (on a scale of 1-10, where 1 is least helpful and 10 is most helpful)
  • Overall, how useful was the proactive chat assistant (e.g., the one that automatically provided suggestions)? (on a scale of 1-10, where 1 is least helpful and 10 is most helpful)
  • Briefly explain the reasoning for your ratings above. (free response)
  • What kind of suggestions did you find most helpful? (free response)
  • In what ways could the proactive chat assistant suggestions be improved?
  • In what ways could the proactive chat assistant interface be improved?
Appendix B Task Design
B.1 Tasks
We consider a total of 4 tasks in our study, where each task comprises multiple sub-tasks.
Task 1: Storefront
You are a freelance software engineer hired to design the backend of an online store. To get started, we have provided starter code for the Store class.
Your goal is to add more functionality to the online store to make it more complete. Please complete Sub-Task 1 first, then you can complete Sub-Tasks 2 and 3 in any order.
Sub-Task 1: major additions - write the Order and Product class, which should work with the current code in the Store class
Sub-Task 2: add the following functionality to the Store class - write function ‘apply_discount_to_order(self, order_id, discount)‘ - write function ‘check_order_status(self, order_id)‘
Sub-Task 3: create one additional feature that you think might be good to have
Make sure to write test cases to demonstrate that the online store works as intended.
You can only submit your code once for this task. Please only submit your code after you have completed as many sub-tasks as you can.
Task 2: To-do List
You are an application developer working on the backend of a to-do list app. You are working with some existing code your colleague left off. Each task takes in a ‘description‘ (string), ‘category‘ (string), ‘priority‘ (string, ‘"High"‘, ‘"Medium"‘, or ‘"Low"‘), and optional ‘date_due‘ (datetime object).
Your goal is to add more functionality to the to-do list to make it more complete. You can complete the subtasks in any order.
Sub-Task 1: minor fixes - when a task is overdue, change the ‘__str__‘ function of the task to reflect the task is overdue by adding ‘(OVERDUE)‘ after the due date - when adding a new task, do not allow the user to add it if the due date is in the past - there seems to be a bug in the code, please find and fix it - the efficiency of editing a task can be improved, please edit this functionality
Sub-Task 2: add features - modify the ‘list_all(self)‘ function by adding an argument ‘show_completed‘ which prints only unfinished tasks when set to False - add a function ‘list_by_priority(self)‘ to ToDoList which prints the task in order of priority, from high to low
Make sure to write test cases to demonstrate that the to-do list works as intended.
You can only submit your code once for this task. Please only submit your code after you have completed as many sub-tasks as you can.
Task 3: Sales analysis
You are a data analyst working for a global retailer. You are working with product sales data from the past year.
The sales data is a 4-dimensional numpy array with shape (3, 12, 10, 100) and the following structure:
- Axis 0: Different regions
- Axis 1: Different months
- Axis 2: Different stores
- Axis 3: Product sales figures
Given the described sales data, you need to use ONLY numpy packages to complete the following sub-tasks:
1. write function ‘total_sales_per_region(data)‘ that calculates total sales number for each region
2. write function ‘cumulative_sales(data)‘ that computes the cumulative sales over time for each product across all regions and stores
3. write function ‘top_products_by_sales(data, k)‘ that computes top k best selling product id for each month across all regions and stores
4. write function ‘temporal_correlation(data)‘ that calculates pairwise temporal correlations over time for each product
Your goal is to complete the above sub-tasks. Make sure to write test cases to demonstrate that the code works as intended.
You can only submit your code once for this task. Please only submit your code after you have completed as many sub-tasks as you can.
Task 4: Weather trends
You are a data analyst working for a meteorological department. Your task is to analyze temperature data for a given month to gain insights into weather patterns and trends.
Given temperature data for a month, you need to use ONLY numpy packages to complete the following sub-tasks:
1. fill out ‘classify_temps(data)‘ which classifies each day’s temperature into categories such as ’Freezing’, ’Moderate’, and ’Hot’.
2. write function ‘clip_temps(data)‘ which clips any extreme temperature values to ensure they fall within -10 and 40.
3. write function ‘compute_moving_avg(data, window_size)‘ which calculates the moving average of temperatures over a specified window_size (e.g., 7 days).
4. write function ‘compute_weekly_avg(data)‘ which calculates weekly average temperatures.
Your goal is to complete the above sub-tasks. The provided code is incomplete and may not be fully correct.
Make sure to write test cases to demonstrate that the code works as intended.
You can only submit your code once for this task. Please only submit your code after you have completed as many sub-tasks as you can.
B.2 Task organization
Since each participant experienced both the baseline and proactive conditions, we distributed the 4 tasks in a symmetrical manner. This means that if the participant started with a system-building task (Task 1 or 2), they would see the other system-building task when the proactive assistant was switched on or off. This scheme ensures that we have a relatively similar number of each task type in both proactive and baseline conditions.
Appendix C Prompts
Debugging prompt
[insert code]
I need help identifying and fixing the issues in the code. Here are some relevant details:
stdout: [insert err.stdout]
stderr: [insert err.stderr]
Suggest 1 to 2 solutions to fix the issues.
Use the following format for your response: 1. short title explanation on what went wrong and how to fix it “‘ code snippet addressing the problem “‘
Preview prompt
We have the following ORIGINAL code:
[insert code]
A suggestion has been provided:
[insert suggestion]
INSTRUCTION: - If the suggestion contains no code or potential modifications to the code, return 0 and nothing else. Only zero - If the suggestions contain code, integrate the suggestion into the ORIGINAL code. Make minimal changes to the ORIGINAL code. Return the changed code only with nothing else.
Suggestion generation prompt
We use the following system prompt:
You are a helpful programming assistant. You can predict what the user wants to do next and provide suggestions. You never make mistakes in your code. Consider things such as explaining existing code, brainstorming new ideas, providing syntax hints and links to external documentations, identifying and fixing bugs, adding unit tests, completing unfinished code, improving code efficiency and modularity, etc. Do not include suggestions only adding docstrings. Do not include suggestions involving packages outside of vanilla python and numpy.
We use the following main prompt:
[insert code]
Provide suggestions based on the above code context. Include ‘function name‘ in suggestion title if it is mentioned in the suggestion.
Use one of the following formats depending on suggestion type, provide 3 suggestions:
1. suggestion type: short title explaining provided code or brainstorming high-level ideas
OR
1. suggestion type: short title “‘ one or more suggested code snippets, do not include ‘python‘ “‘ clear and detailed explanation for each code snippet in bullet point format. If code is very straightforward then don’t explain
"suggestion type" can be one of (Code Explanation, Code Improvement, Brainstorming Idea, Testing, Bug Fix, Syntax Hint) or something else. If there are multiple functions or classes in the code, provide references to the specific function or class that the suggestion pertains to in the explanation.